# Iwin Transformer: Hierarchical Vision Transformer using Interleaved Windows

### v1

@@ Abstract
Index Terms:
Iwin Transformer, interleaved window attention, position-embedding-free.
@@ I Introduction
Vision Transformers (ViTs) [1] have have fundamentally transformed computer vision by borrowing the transformer architecture from natural language models [2]. Unlike Convolutional Neural Networks (CNNs) [3], which rely on local receptive fields to capture image features, ViTs leverage self-attention mechanisms to get global dependencies, demonstrating remarkable performance on vision tasks. However, its quadratic computational complexity $\mathcal{O}(N^{2})$ with respect to input sequence length $N$ presents significant scalability challenges, particularly for high-resolution image processing applications that are increasingly common in computer vision.
@@
To tackle the challenge of quadratic complexity in Vision Transformers (ViTs) and enhance their efficiency while maintaining performance, various approaches have been proposed. Hierarchical Designs such as PVT [4] and Twins [5] utilize multi-scale feature pyramids to progressively reduce spatial dimensions. Hybrid CNN-Transformer Architectures like ConViT [6] and CoAtNet [7] combine convolutional operations with self-attention to leverage the strengths of both paradigms. Efficient Token Fusion strategies such as TokenLearner [8] dynamically aggregate tokens to reduce sequence length, while Sparse Attention Patterns exemplified by Reformer [9] utilize locality-sensitive hashing to attend only to relevant tokens. Additionally, efficient implementations like Performer [10] approximate attention through kernel methods to achieve linear complexity. Diverse strategies are employed to mitigate the computational demands of vision transformers.
@@
One of the most promising approaches to tackling these challenges is the Swin Transformer [11], which introduces a hierarchical architecture with shifted window-based self-attention. By constraining attention computation within local windows and enabling cross-window connection through a window-shifting mechanism, Swin Transformer successfully reduces the quadratic complexity to linear complexity with respect to image size. This elegant design maintains the model’s capability to capture both local and global dependencies while significantly improving computational efficiency. Moreover, Swin Transformer adopts a hierarchical structure that progressively merges image patches in deeper layers, generating multi-scale feature maps similar to conventional CNN backbones, which facilitates its application across various vision tasks including object detection and semantic segmentation. The impressive performance of Swin Transformer across benchmarks has established it as a milestone in efficient vision transformer design and demonstrated the viability of window-based attention mechanisms for large-scale vision applications.
@@
Despite its pioneering design and impressive performance, the Swin Transformer exhibits several noteworthy limitations. First, the shifted window mechanism introduces extra computational overhead due to the complex masking operations required during attention computation, complicating implementation and reducing hardware efficiency. Second, Swin’s architecture requires two consecutive transformer blocks: one with regular windows and another with shifted windows to achieve global information exchange, resulting in computational redundancy as certain features are processed multiple times. This two-block requirement poses particular challenges in the era of AI generated content (AIGC), where conditioning information such as text prompts must be injected into the model; there is no obvious optimal placement for cross-attention between text and images within this rigid two-block structure, explaining Swin’s limited adoption in modern text-to-image diffusion models. Furthermore, as acknowledged in Swin Transformer v2 [12], the model faces scalability issues when fine-tuned for higher-resolution inputs. The bi-cubic interpolation of relative position encodings for larger windows leads to significant performance degradation, necessitating the introduction of complex alternatives such as log-spaced continuous position bias (Log-CPB). This reliance on sophisticated position encoding schemes ultimately hinders the model’s scaling capabilities and broader applicability.
@@ II Related Works
Inspired by the success of Vision Transformers (ViTs) [1], Transformer architectures have drawn significant attention in computer vision research [13, 14, 15, 16, 17, 18]. While they all encounter one common problem, which is the heavy computational overhead of quadratic complexity. Various approaches have been proposed to enhance the efficiency of transformer structure while mataining good performance. Some works [6, 7, 19] integrate CNNs and Transformers to leverage the advantages of both structures. Other approaches  [4, 11, 20] focus on modifying the structure of ViTs to better suit vision tasks. The following subsections briefly review these related works categorized by their methodologies.
@@ II-A Linear and Sparse Attention
Self-attention operations introduce quadratic computational complexity with respect to sequence length, presenting significant challenges for high-resolution visual inputs. Linformer [21] achieved linear complexity through low-rank factorization of attention matrices, decomposing the $N\times N$ attention matrix into two smaller matrices. Performer [10] introduced Fast Attention Via positive Orthogonal Random features (FAVOR+), using random feature maps to approximate the attention kernel. Luna [22] proposed linear unified nested attention by introducing a set of fixed-length projected embeddings that serve as an intermediate representation for attention computation. BigBird [23] combined random, window, and global attention patterns to maintain linear complexity. Longformer [24] employed dilated sliding window attention with select global attention tokens. Sparse Transformer [25] introduced factorized attention patterns that reduce complexity through structured sparsity. [26, 27] selectively compute attention for the most relevant token pairs. These approaches collectively demonstrate that full global attention is often unnecessary for effective visual representation learning, enabling more efficient transformer designs without substantial performance degradation.
@@ II-B Hierarchical Vision Transformers
Hierarchical vision transformers adopt multi-scale feature representations to enhance computational efficiency while maintaining modeling capabilities. Pyramid Vision Transformer (PVT) [4] introduces a progressive shrinking pyramid structure that reduces sequence length at deeper layers through spatial-reduction attention. Swin Transformer [11] presents a hierarchical architecture with shifted windows, restricting self-attention computation to local windows and establishing cross-window connections through window shifting between layers. This design reduces computational complexity from quadratic to linear with respect to image size. MViT [20] employs pooling attention that progressively expands channel capacity while reducing spatial resolution. Twins combines locally-grouped self-attention with global sub-sampled attention to balance local and global interactions efficiently. CSWin [28] utilizes cross-shaped window self-attention to capture horizontal and vertical dependencies separately. These hierarchical designs have proven particularly effective for dense prediction tasks like object detection and semantic segmentation, where multi-scale feature representations are essential.
@@ II-C Hybrid CNN-Transformer Architectures
Hybrid architectures combine convolutional operations with self-attention to leverage the strengths of both paradigms. ConViT [6] incorporates a gated positional self-attention mechanism that transitions smoothly from convolution to transformer behavior. CoAtNet [7] unifies depthwise convolutions with self-attention in a relative attention framework, employing convolutions in earlier stages and self-attention in later stages. MobileViT [19] integrates the local processing of convolutions with the global processing of transformers for lightweight models. LocalViT [29] enhances vision transformers with depth-wise convolutions, introducing locality to self-attention layers. These hybrid approaches typically achieve better parameter efficiency and performance on smaller datasets compared to pure transformer architectures, while maintaining the global modeling capacity essential for complex vision tasks.
@@ II-D Dynamic Computation Strategies
Dynamic computation strategies adjust computational resources based on input complexity. DynamicViT [30] introduces a token sparsification framework that progressively prunes redundant tokens based on their importance scores, reducing sequence length as the network deepens. A-ViT [31] employs adaptive computation depth, allowing different tokens to exit the network at different layers based on their complexity. Token Merging (ToMe) [32] dynamically merges similar tokens throughout the network, preserving information while reducing sequence length. Adaptive Token Sampling [33] introduces learnable modules that sample important tokens based on the input, preserving critical information while reducing computation. These approaches enable more efficient processing by allocating computational resources where they are most needed, making them particularly suitable for real-time applications with varying input complexities.
@@ II-E Difference With Previous Works
Unlike Swin Transformer [11], which requires two consecutive blocks with regular and shifted window patterns to establish cross-window connections, Iwin Transformer achieves global information exchange within a single block through a synergistic combination of interleaved window attention and depthwise separable convolution. Any token can interact with others through at intermediary connection, similar to a flat organization where individuals can reach others without going through multiple people. Another benefit is the module composed of interleaved window attention and depthwise separable convolution can seamlessly replace the standard attention module in generative models without affecting subsequent cross-attention operations with text conditions. Swin can not do it due to its two-block dependency.
In contrast to previoyus hybrid CNN-Transformer architectures like LocalViT [29] and MobileViT [19], which typically use convolutions merely to capture local features as the motivation. In Iwin Transformer, depthwise separable convolution and interleaved window attention are interdependent components forming a cohesive information processing unit. Depthwise separable convolutions establish connections between certain tokens that are not established through interleaved window attention. Another advantage is that, because convolutions naturally carry position information, Iwin no longer requires explicit position encoding. This makes Iwin a position-embedding-free transformer, which allows models trained on low resolution to be easily fine-tuned to high resolution while maintaining performance.
@@ III-A Overall Architecture
An overview of the Iwin Transformer architecture is presented in Figure 7. The detailed configuration is shown in Table IIIIwin Transformer follows a hierarchical architecture similar to Swin Tansformer [11], progressively reduces spatial resolution while expanding channel dimensions across four stages. Given an input image, Iwin first splits it into non-overlapping patches by a patch splitting module. Each patch is treated as a token or feature, then architecture processes features through stages $\{S_{1},S_{2},S_{3},S_{4}\}$ with resolutions $\{H/4,H/8,H/16,H/32\}$ and channels $\{C,2C,4C,8C\}$ respectively. Although pyramid structures were used in this study, the core modules of Iwin can be used in flat structures.
@@ III-B1 Rearrange
For a token originally at position $(i,j)$, it will be assigned to the window represented by $(window\_row,window\_col)$:
@@
Note that $i\bmod H_{g}$ ranges from 00 to $H_{g}-1$, which exactly corresponds to the row index of the window in the grid of windows. Similarly, $window\_col=j\bmod W_{g}$ ranges from 00 to $W_{g}-1$, corresponding to the column index of the window.
@@ III-B2 Self-Attention
Within each window, we apply the standard self-attention mechanism [2]:
@@ III-B3 Restore
III-C Depthwise Separable Convolution
Depthwise Separable Convolution (DWConv) [34] is used to help build missing relationships of certain tokens which are not in the same attention window and provide implicit positional information by the way.
@@ III-D Downsampling Layer
Iwin Transformer employs standard convolution to progressively reduce spatial resolution and increase channel dimensions, following the hierarchical design principle  [35] common in vision backbones.
@@ III-E Iwin Transformer Block
As shown in Figure  3a, the Iwin Transformer block consists of a unified module integrating Interleaved Window Multi-Head Self-Attention (IW-MSA) and Depthwise Separable Convolution (DWConv) modules in parallel, followed by a two-layer MLP with GELU [36] activation between layers. A LayerNorm layer precedes each unified module and MLP, with residual connections after each. The forward pass of the Iwin Transformer block can be formulated as:
@@ III-F Architecture Variants
Referring to the Swin Transformer [11], we build Iwin-T, Iwin-S, Iwin-B, and Iwin-L with identical network depth and width as Swin for fair comparison. And it is also supported by our ablation studies on Iwin-T, which revealed that the layer configuration $\{2,2,6,2\}$ achieves highest accuracy (see Table VII). Therefore, the subsequent Iwin-S, Iwin-B, and Iwin-L all follow the same settings as Swin. For input resolutions of 224, 384, 512, and 1024, window sizes are 7, 12, 16, and 16 respectively. Table I shows the model size, computational complexity (FLOPs), and performance of different variants on ImageNet.
@@ III-G Global Information Exchange
At first, we followed the rule $KM\geq\max(H,W)$, assigning different convolution kernel sizes to each stage. However, ablation experiments (see in Table VII) on Iwin-T showed that kernel sizes of 7, 5, and 3 for stages 1, 2, and 3 yielded the worst performance, with lowest accuracy and slowest training speed, compared to using a consistent kernel size of 7, 5, or 3 across stages. This indicates consistent kernel sizes lead to faster training and better optimization. This aligns with observations in [37] that balanced network outperform theoretically optimal but imbalanced configurations. We think that as the network deepens and downsampling increases the effective receptive field (ERF) [38], resulting in $K_{ERF}\cdot M\geq\max(H,W)$
@@ IV Experiments
We conduct experiments on ImageNet-1K image classification [39], COCO object detection [40], ADE20K semantic segmentation [41], Kinetics-400 [42] video recognization and class-conditional image generation. In the following, we first compare the proposed Iwin Transformer with the previous state-of-the-arts. Then, we ablate the important design elements of Iwin Transformer.
@@ IV-A Image Classification on ImageNet-1K
For image classification, we benchmark the proposed Iwin Transformer on ImageNet-1K [39], which contains 1.28M training images and 50K validation images from 1,000 classes. Experimental settings closely follow those of Swin Transformer [11]. We report top-1 accuracy on a single crop under two scenarios: training from scratch on ImageNet-1K and pre-training on ImageNet-22K, which contains 14.2M images and 22K classes, followed by fine-tuning on ImageNet-1K.
From scratch on ImageNet-1K. We use an AdamW optimizer [43] for 300 epochs with a 20-epoch linear warmup and a cosine learning rate decay. We set the batch size to 512, weight decay to 0.05, and use most of the augmentation strategies from DeiT [44]. The initial learning rate is set to 0.0005, and the drop path rates are 0.2, 0.3, and 0.5 for Iwin-T, Iwin-S, and Iwin-B, respectively.
Pre-training on ImageNet-22K. Training lasts for 90 epochs with a 5-epoch warmup. We use a batch size of 4096, an initial learning rate of $41.25e-$ 4, and a weight decay of 0.05. The pre-trained models are then fine-tuned on ImageNet-1K for 10 30 epochs at a $224\times 224$ resolution, using a batch size of 1024, a constant learning rate of $052e-$ 05, and a weight decay of $81e-$ 8.
@@
During this cross-resolution fine-tuning, we adjust the window size to match the input resolution—7 for 224, 12 for 384, and 16 for 512 and 1024—while keeping the architectures unchanged. The fine-tuning process runs for 30 epochs with a constant learning rate of $052e-$ 05, a weight decay of $81e-$ 8, and a warm-up of 5 to 10 epochs.

### v2

@@ I Introduction
Vision Transformers (ViTs) [dosovitskiy2020image] have have fundamentally transformed computer vision by borrowing the transformer architecture from natural language models [vaswani2017attention]. Unlike Convolutional Neural Networks (CNNs) [krizhevsky2012imagenet], which rely on local receptive fields to capture image features, ViTs leverage self-attention mechanisms to get global dependencies, demonstrating remarkable performance on vision tasks. However, its quadratic computational complexity $\mathcal{O}(N^{2})$ with respect to input sequence length $N$ presents significant scalability challenges, particularly for high-resolution image processing applications that are increasingly common in computer vision.
@@
To tackle the challenge of quadratic complexity in Vision Transformers (ViTs) and enhance their efficiency while maintaining performance, various approaches have been proposed. Hierarchical Designs such as PVT [wang2021pyramid] and Twins [chu2021twins] utilize multi-scale feature pyramids to progressively reduce spatial dimensions. Hybrid CNN-Transformer Architectures like ConViT [d2021convit] and CoAtNet [dai2021coatnet] combine convolutional operations with self-attention to leverage the strengths of both paradigms. Efficient Token Fusion strategies such as TokenLearner [ryoo2021tokenlearner] dynamically aggregate tokens to reduce sequence length, while Sparse Attention Patterns exemplified by Reformer [kitaev2020reformer] utilize locality-sensitive hashing to attend only to relevant tokens. Additionally, efficient implementations like Performer [choromanski2020rethinking] approximate attention through kernel methods to achieve linear complexity. Diverse strategies are employed to mitigate the computational demands of vision transformers.
@@
One of the most promising approaches to tackling these challenges is the Swin Transformer [liu2021swin], which introduces a hierarchical architecture with shifted window-based self-attention. By constraining attention computation within local windows and enabling cross-window connection through a window-shifting mechanism, Swin Transformer successfully reduces the quadratic complexity to linear complexity with respect to image size. This elegant design maintains the model’s capability to capture both local and global dependencies while significantly improving computational efficiency. Moreover, Swin Transformer adopts a hierarchical structure that progressively merges image patches in deeper layers, generating multi-scale feature maps similar to conventional CNN backbones, which facilitates its application across various vision tasks including object detection and semantic segmentation. The impressive performance of Swin Transformer across benchmarks has established it as a milestone in efficient vision transformer design and demonstrated the viability of window-based attention mechanisms for large-scale vision applications.
@@
Despite its pioneering design and impressive performance, the Swin Transformer exhibits several noteworthy limitations. First, the shifted window mechanism introduces extra computational overhead due to the complex masking operations required during attention computation, complicating implementation and reducing hardware efficiency. Second, Swin’s architecture requires two consecutive transformer blocks: one with regular windows and another with shifted windows to achieve global information exchange, resulting in computational redundancy as certain features are processed multiple times. This two-block requirement poses particular challenges in the era of AI generated content (AIGC), where conditioning information such as text prompts must be injected into the model; there is no obvious optimal placement for cross-attention between text and images within this rigid two-block structure, explaining Swin’s limited adoption in modern text-to-image diffusion models. Furthermore, as acknowledged in Swin Transformer v2 [liu2022swin], the model faces scalability issues when fine-tuned for higher-resolution inputs. The bi-cubic interpolation of relative position encodings for larger windows leads to significant performance degradation, necessitating the introduction of complex alternatives such as log-spaced continuous position bias (Log-CPB). This reliance on sophisticated position encoding schemes ultimately hinders the model’s scaling capabilities and broader applicability.
@@ II Related Works
Inspired by the success of Vision Transformers (ViTs) [dosovitskiy2020image], Transformer architectures have drawn significant attention in computer vision research [touvron2021training, carion2020end, zheng2021rethinking, arnab2021vivit, zhu2020deformable, peebles2023scalable]. While they all encounter one common problem, which is the heavy computational overhead of quadratic complexity. Various approaches have been proposed to enhance the efficiency of transformer structure while mataining good performance. Some works [d2021convit, dai2021coatnet, mehta2021mobilevit] integrate CNNs and Transformers to leverage the advantages of both structures. Other approaches  [wang2021pyramid, liu2021swin, fan2021multiscale] focus on modifying the structure of ViTs to better suit vision tasks. The following subsections briefly review these related works categorized by their methodologies.
@@ II-A Linear and Sparse Attention
Self-attention operations introduce quadratic computational complexity with respect to sequence length, presenting significant challenges for high-resolution visual inputs. Linformer [wang2020linformer] achieved linear complexity through low-rank factorization of attention matrices, decomposing the $N\times N$ attention matrix into two smaller matrices. Performer [choromanski2020rethinking] introduced Fast Attention Via positive Orthogonal Random features (FAVOR+), using random feature maps to approximate the attention kernel. Luna [ma2021luna] proposed linear unified nested attention by introducing a set of fixed-length projected embeddings that serve as an intermediate representation for attention computation. BigBird [zaheer2020big] combined random, window, and global attention patterns to maintain linear complexity. Longformer [beltagy2020longformer] employed dilated sliding window attention with select global attention tokens. Sparse Transformer [child2019generating] introduced factorized attention patterns that reduce complexity through structured sparsity. [tay2020sparse, roy2021efficient] selectively compute attention for the most relevant token pairs. These approaches collectively demonstrate that full global attention is often unnecessary for effective visual representation learning, enabling more efficient transformer designs without substantial performance degradation.
@@ II-B Hierarchical Vision Transformers
Hierarchical vision transformers adopt multi-scale feature representations to enhance computational efficiency while maintaining modeling capabilities. Pyramid Vision Transformer (PVT) [wang2021pyramid] introduces a progressive shrinking pyramid structure that reduces sequence length at deeper layers through spatial-reduction attention. Swin Transformer [liu2021swin] presents a hierarchical architecture with shifted windows, restricting self-attention computation to local windows and establishing cross-window connections through window shifting between layers. This design reduces computational complexity from quadratic to linear with respect to image size. MViT [fan2021multiscale] employs pooling attention that progressively expands channel capacity while reducing spatial resolution. Twins combines locally-grouped self-attention with global sub-sampled attention to balance local and global interactions efficiently. CSWin [dong2022cswin] utilizes cross-shaped window self-attention to capture horizontal and vertical dependencies separately. These hierarchical designs have proven particularly effective for dense prediction tasks like object detection and semantic segmentation, where multi-scale feature representations are essential.
@@ II-C Hybrid CNN-Transformer Architectures
Hybrid architectures combine convolutional operations with self-attention to leverage the strengths of both paradigms. ConViT [d2021convit] incorporates a gated positional self-attention mechanism that transitions smoothly from convolution to transformer behavior. CoAtNet [dai2021coatnet] unifies depthwise convolutions with self-attention in a relative attention framework, employing convolutions in earlier stages and self-attention in later stages. MobileViT [mehta2021mobilevit] integrates the local processing of convolutions with the global processing of transformers for lightweight models. LocalViT [li2021localvit] enhances vision transformers with depth-wise convolutions, introducing locality to self-attention layers. These hybrid approaches typically achieve better parameter efficiency and performance on smaller datasets compared to pure transformer architectures, while maintaining the global modeling capacity essential for complex vision tasks.
@@ II-D Dynamic Computation Strategies
Dynamic computation strategies adjust computational resources based on input complexity. DynamicViT [rao2021dynamicvit] introduces a token sparsification framework that progressively prunes redundant tokens based on their importance scores, reducing sequence length as the network deepens. A-ViT [yin2022vit] employs adaptive computation depth, allowing different tokens to exit the network at different layers based on their complexity. Token Merging (ToMe) [bolya2022token] dynamically merges similar tokens throughout the network, preserving information while reducing sequence length. Adaptive Token Sampling [fayyaz2022adaptive] introduces learnable modules that sample important tokens based on the input, preserving critical information while reducing computation. These approaches enable more efficient processing by allocating computational resources where they are most needed, making them particularly suitable for real-time applications with varying input complexities.
@@ II-E Difference With Previous Works
Unlike Swin Transformer [liu2021swin], which requires two consecutive blocks with regular and shifted window patterns to establish cross-window connections, Iwin Transformer achieves global information exchange within a single block through a synergistic combination of interleaved window attention and depthwise separable convolution. Any token can interact with others through at intermediary connection, similar to a flat organization where individuals can reach others without going through multiple people. Another benefit is the module composed of interleaved window attention and depthwise separable convolution can seamlessly replace the standard attention module in generative models without affecting subsequent cross-attention operations with text conditions. Swin can not do it due to its two-block dependency.
In contrast to previoyus hybrid CNN-Transformer architectures like LocalViT [li2021localvit] and MobileViT [mehta2021mobilevit], which typically use convolutions merely to capture local features as the motivation. In Iwin Transformer, depthwise separable convolution and interleaved window attention are interdependent components forming a cohesive information processing unit. Depthwise separable convolutions establish connections between certain tokens that are not established through interleaved window attention. Another advantage is that, because convolutions naturally carry position information, Iwin no longer requires explicit position encoding. This makes Iwin a position-embedding-free transformer, which allows models trained on low resolution to be easily fine-tuned to high resolution while maintaining performance.
@@ III-A Overall Architecture
An overview of the Iwin Transformer architecture is presented in Figure 7. The detailed configuration is shown in Table IIIIwin Transformer follows a hierarchical architecture similar to Swin Tansformer [liu2021swin], progressively reduces spatial resolution while expanding channel dimensions across four stages. Given an input image, Iwin first splits it into non-overlapping patches by a patch splitting module. Each patch is treated as a token or feature, then architecture processes features through stages $\{S_{1},S_{2},S_{3},S_{4}\}$ with resolutions $\{H/4,H/8,H/16,H/32\}$ and channels $\{C,2C,4C,8C\}$ respectively. Although pyramid structures were used in this study, the core modules of Iwin can be used in flat structures.
@@ III-B1 Rearrange
For a token originally at position $(i,j)$, it will be assigned to the window represented by (window_row,window_col)(window\_row,window\_col):
@@
Note that $i\bmod H_{g}$ ranges from 0 to $H_{g}-1$, which exactly corresponds to the row index of the window in the grid of windows. Similarly, $window\_col=j\bmod W_{g}$ ranges from 0 to $W_{g}-1$, corresponding to the column index of the window.
@@ III-B2 Self-Attention
Within each window, we apply the standard self-attention mechanism [vaswani2017attention]:
@@ III-B3 Restore
III-C Depthwise Convolution
Depthwise Convolution (DWConv) [chollet2017xception] is used to help build missing relationships of certain tokens which are not in the same attention window and provide implicit positional information by the way.
@@ III-D Downsampling Layer
Iwin Transformer employs standard convolution to progressively reduce spatial resolution and increase channel dimensions, following the hierarchical design principle  [he2016deep] common in vision backbones.
@@ III-E Iwin Transformer Block
As shown in Figure  3a, the Iwin Transformer block consists of a unified module integrating Interleaved Window Multi-Head Self-Attention (IW-MSA) and Depthwise Separable Convolution (DWConv) modules in parallel, followed by a two-layer MLP with GELU [hendrycks2016gaussian] activation between layers. A LayerNorm layer precedes each unified module and MLP, with residual connections after each. The forward pass of the Iwin Transformer block can be formulated as:
@@ III-F Architecture Variants
Referring to the Swin Transformer [liu2021swin], we build Iwin-T, Iwin-S, Iwin-B, and Iwin-L with identical network depth and width as Swin for fair comparison. And it is also supported by our ablation studies on Iwin-T, which revealed that the layer configuration $\{2,2,6,2\}$ achieves highest accuracy (see Table VII). Therefore, the subsequent Iwin-S, Iwin-B, and Iwin-L all follow the same settings as Swin. For input resolutions of 224, 384, 512, and 1024, window sizes are 7, 12, 16, and 16 respectively. Table I shows the model size, computational complexity (FLOPs), and performance of different variants on ImageNet.
@@ III-G Global Information Exchange
At first, we followed the rule $KM\geq\max(H,W)$, assigning different convolution kernel sizes to each stage. However, ablation experiments (see in Table VII) on Iwin-T showed that kernel sizes of 7, 5, and 3 for stages 1, 2, and 3 yielded the worst performance, with lowest accuracy and slowest training speed, compared to using a consistent kernel size of 7, 5, or 3 across stages. This indicates consistent kernel sizes lead to faster training and better optimization. This aligns with observations in [tan2019efficientnet] that balanced network outperform theoretically optimal but imbalanced configurations. We think that as the network deepens and downsampling increases the effective receptive field (ERF) [luo2016understanding], resulting in $K_{ERF}\cdot M\geq\max(H,W)$ .
@@ IV Experiments
We conduct experiments on ImageNet-1K image classification [deng2009imagenet], COCO object detection [lin2014coco], ADE20K semantic segmentation [zhou2019semantic], Kinetics-400 [kay2017kinetics] video recognization and class-conditional image generation. In the following, we first compare the proposed Iwin Transformer with the previous state-of-the-arts. Then, we ablate the important design elements of Iwin Transformer.
@@ IV-A Image Classification on ImageNet-1K
For image classification, we benchmark the proposed Iwin Transformer on ImageNet-1K [deng2009imagenet], which contains 1.28M training images and 50K validation images from 1,000 classes. Experimental settings closely follow those of Swin Transformer [liu2021swin]. We report top-1 accuracy on a single crop under two scenarios: training from scratch on ImageNet-1K and pre-training on ImageNet-22K, which contains 14.2M images and 22K classes, followed by fine-tuning on ImageNet-1K.
From scratch on ImageNet-1K. We use an AdamW optimizer [kingma2017adam] for 300 epochs with a 20-epoch linear warmup and a cosine learning rate decay. We set the batch size to 512, weight decay to 0.05, and use most of the augmentation strategies from DeiT [touvron2020deit]. The initial learning rate is set to 0.0005, and the drop path rates are 0.2, 0.3, and 0.5 for Iwin-T, Iwin-S, and Iwin-B, respectively.
Pre-training on ImageNet-22K. Training lasts for 90 epochs with a 5-epoch warmup. We use a batch size of 4096, an initial learning rate of 1.25e41.25e-4, and a weight decay of 0.05. The pre-trained models are then fine-tuned on ImageNet-1K for 10 30 epochs at a $224\times 224$ resolution, using a batch size of 1024, a constant learning rate of 2e052e-05, and a weight decay of 1e81e-8.
@@
During this cross-resolution fine-tuning, we adjust the window size to match the input resolution—7 for 224, 12 for 384, and 16 for 512 and 1024—while keeping the architectures unchanged. The fine-tuning process runs for 30 epochs with a constant learning rate of 2e052e-05, a weight decay of 1e81e-8, and a warm-up of 5 to 10 epochs.